OpenStudy (anonymous):

Help Determine which of the following are solutions to the equation below. x^2 = 14 A. 14 B. 7 C. ± square root of 14 D. ± square root of 7 I don't know what "±" means, could someone tell me?

OpenStudy (anonymous):

\[\left( -\sqrt{14} \right)^{2}=14\] \[\left( \sqrt{14} \right)^{2}=14\] \[\pm~ means~ value~ can~ be + ~or~ -\] 1 hope you got it.